The amount of annual snowfall in a certain mountain range is normally distributed with a mean of 72 inches and a standard deviation of 11 inches. What is the probability that the annual snowfall for 1 randomly picked year will exceed 80 inches? Round your answer to 3 decimal places.
Answer #1

Solution :

Given that ,

P(x > 80) = 1 - P(x < 80)

= 1 - P[(x - ) / < (80 - 72) / 11)

= 1 - P(z < 0.73)

= 1 - 0.7673

= 0.233

Probability = 0.233